# Smeared phase transition in the dissipative random quantum Ashkin-Teller model

### v1

@@ V Strong-disorder renormalization group
We now present the SDRG decimation procedure relevant for the $\epsilon_{I}>1$ regime in the presence of dissipation. We follow the methods of the dissipationless case in Ref. (2014_ashkinteller) and include dissipation using the strategy of Ref. (smeared_hoyos_2) which is the following. The renormalization-group decimations are of two types: global and local, and are performed simultaneously when lowering the cutoff energy scale of the system. The global decimation concerns the integration of the fastest bath oscillators and is related to the adiabatic renormalization (see Sec. IV). The local decimation concerns the coarse-grainning of a high-energy localized mode in a disordered system and, thus, is identical to the dissipationless case.
@@
where $\Omega_{I}$ is the bare $\Omega$. Notice that this is the global decimation procedure mentioned above. In contrast, the remaining steps are local decimations and are identical to the dissipationless case, except when it comes to renormalize the dissipation parameter $\alpha_{i}$.
(ii) Let the two largest energy scales be $g_{i}$ and $K_{i}$. Thus, the unperturbed Hamiltonian is $H_{0}=-K_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}-g{}_{i}\eta_{i}^{x}$ and all other terms in the Hamiltonian (4) containing $\boldsymbol{\sigma}_{i}$, $\boldsymbol{\sigma}_{i+1}$, and $\boldsymbol{\eta}_{i}$ are treated perturbatively up to second order. As a result, the new effective Hamiltonian has the same form as the Hamiltonian (4) but with one fewer site and two fewer spin-1/2 degrees of freedom. The degrees of freedom $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\sigma}_{i+1}$ are clustered together and replaced by an effective spin-1/2 degree of freedom $\tilde{\boldsymbol{\sigma}}$ upon which the effective field is $\tilde{h}$. In addition, $\boldsymbol{\eta}_{i}$ becomes locked in the $x$-direction and does not contribute to the $z$-magnetization. The effective coupling between the neighboring $\boldsymbol{\eta}$’s is $\tilde{J}$. This is sketched in Fig. 2(a), and the renormalized couplings are
$\tilde{J}=\frac{2J_{i-1}J_{i}}{\tilde{g}_{i}},\quad\tilde{h}=\frac{2h_{i}h_{i+1}}{K_{i}}.$ (20)
Now, we need to compute how the dissipation strength $\alpha$ changes under this decimation. Recalling that
@@
and that the doubly degenerated ground state is $\left\{\left|\rightarrow_{i};\uparrow_{i}\uparrow_{i+1}\right\rangle,\left|\rightarrow_{i};\downarrow_{i}\downarrow_{i+1}\right\rangle\right\}$. Then, projecting $H_{\text{SB}}$ onto these states, we find that
@@
This is simply because $\left\langle\eta_{i}^{z}\right\rangle=\left\langle\eta_{i}^{z}\sigma_{i}^{z}\right\rangle=0$ and $\left\langle\sigma_{i+1}^{z}\right\rangle=\tilde{\sigma}^{z}$. Actually, this could be obtained in another fashion. Since $g_{i}$ is the largest energy scale, it means that the associated degree of freedom is actually already (or nearly) decoupled from the associated bath via successive adiabatic renormalizations (see Sec. IV). Thus, only the last term in (21) needs to be projected onto the ground-state manifold. Importantly, there is no change in the spectral function of the remaining baths, and, therefore, $\alpha_{i+1}$ remains unchanged.
Figure 2: Decimation procedure when the two largest local energy scales in the system are (a) $g_{2}$ and $K_{2}$, (b) $g_{2}$ and $h_{2}$, and (c) $K_{2}$ and $J_{2}$ [in the main text, these correspond to the decimation cases (ii), (iii) and (iv), respectively]. The couplings to the oscillator baths also renormalized and are not illustrated here.
(iii) Let the two largest energy scales be $g_{i}$ and $h_{i}$. In this case, all tunneling frequencies have fully (or nearly) converged to their fixed-point values $g^{*}$ and $h^{*}$ in the adiabatic renormalization. The SDRG procedure is, thus, the same as the dissipationless case. The unperturbed Hamiltonian to be considered is $H_{0}=-h_{i}\sigma_{i}^{x}\left(1+\eta_{i}^{x}\right)-g_{i}\eta_{i}^{x}$. Projecting the spins $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\eta}_{i}$ onto the ground state of $H_{0}$ means that both spins are going to be decimated. The renormalized system is that illustrated in Fig. 2(b) and the renormalized quantities are
$\tilde{K}=\frac{K_{i-1}K_{i}}{2h_{i}},\quad\tilde{J}=\frac{J_{i-1}J_{i}}{g_{i}+h_{i}}.$ (23)
@@
(iv) The final case to consider is when $K_{i}$ and $J_{i}$ are the largest energy scales. The unperturbed Hamiltonian is $H_{0}=-K_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}-J_{i}\eta_{i}^{z}\eta_{i+1}^{z}-J_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}\eta_{i}^{z}\eta_{i+1}^{z}$. Projecting $\boldsymbol{\sigma}_{i}$, $\boldsymbol{\sigma}_{i+1}$, $\boldsymbol{\eta}_{i}$, and $\boldsymbol{\eta}_{i+1}$ onto the ground state means that those spins are always aligned and, therefore, can be represented by two effective degrees of freedom $\tilde{\boldsymbol{\sigma}}$ and $\tilde{\boldsymbol{\eta}}$. Fusing them into clusters is illustrated in Fig. 2(c) where the renormalized quantities are
$\tilde{g}=\frac{g_{i}g_{i+1}}{2J_{i}},\quad\tilde{h}=\frac{h_{i}h_{i+1}}{K_{i}+J_{i}}.$ (24)
For this step, we need to take a closer look on what happens to the oscillator baths. The system-bath Hamiltonian is
@@
and needs to be projected onto the 4-fold-degenerate ground state of $H_{0}$, which is $\left\{\left|\eta;\sigma\right\rangle\right\}$ (where $\eta_{j}^{z}\left|\eta;\sigma\right\rangle=\eta\left|\eta;\sigma\right\rangle$, $\sigma_{j}^{z}\left|\eta;\sigma\right\rangle=\sigma\left|\eta;\sigma\right\rangle$, with $j=i,i+1$ and $\eta,\sigma=\pm 1$ representing the effective spins). Since $H_{\text{SB}}$ is diagonal in all four states, the projection can be easily performed, yielding the following effective Hamiltonian
@@
where the sum in the index $j$ run from $i$ to $i+1$. Consequently, the spectral density of the local renormalized bath is
@@
It is thus interesting to notice that $\tilde{\alpha}$, increases under the RG flow only when FM-like clusterings take place, i.e., only when $J_{i}$ is decimated. Moreover, from Eq. (28), it becomes clear that the effective dissipation strength increases as the volume of the FM cluster, just like in the case of a direct transition between the PM and FM phases studied in Refs. (smeared_hoyos_1; smeared_hoyos_2). In addition, it is possible to relate the dissipation strength with the effective magnetic moment as both increases with the cluster volume. It is a convenient relation when analyzing the SDRG flow equations for the distributions of couplings, fields and dissipation parameters (smeared_hoyos_2).
VI Phase diagram
@@
We are now able to describe the phase diagram of our model Hamiltonian (5) sketched in Fig. 1. A phase diagram is obtained after analyzing the corresponding strong-disorder renormalization-group flow of the Hamiltonian parameters. A careful reader may note that the decimation cases (i)–(iv) described in Sec. V do not cover all possible decimations. However, they do cover all decimations necessary to analyze the PM-PROD and PROD-FM transitions as shown in Ref. (2014_ashkinteller). To review the undamped case, we start noticing that the PM-PROD transition happens from the competition between the transverse field $h$ and the 4-spin coupling $K$ [see the first sum in the Hamiltonian (4), the other terms are just bystanders]. The relevant decimations are, thus, those described in cases (ii) and (iii) in the previous Section. Comparing the corresponding renormalized values $\tilde{h}$ and $\tilde{K}$ in Eqs. (20) and (23), the transition line happens when the typical values are $2h_{\text{typ}}=K_{\text{typ}}$. Thus, the transition line in the undamped case is at $\delta\approx\epsilon_{I}/2$. For the PROD-FM transition, the competition is between the 2-spin couplings $J$ and $g$ [see the second sum in the Hamiltonian (4)]. The relevant decimations are thus the cases (ii) and (iv). Following the same reasoning as before, the transition line is at $\delta\approx 2/\epsilon_{I}.$
@@
Let us now analyze the effects of weak Ohmic dissipation $\alpha\ll 1$ and $s=1$ in Eq. (8)] in the PM-PROD transition. Since $\alpha\ll 1$ and does not renormalize, the decimation of fast oscillators has little effect on the renormalized values of $h$’s [see Eq. (18)]. Therefore, Ohmic dissipation is a irrelevant perturbation to the PM-PROD quantum phase transition. Its only effect is a slight change of the critical line. This same conclusion applies to the case of super-Ohmic dissipation ($s>1$) which is qualitatively weaker when compared to Ohmic dissipation (leggett_spinboson). Very likely, the same conclusion may also apply to the case of sub-Ohmic dissipation ($s<1$).
@@
The effects of dissipation is much more dramatic in the PROD-FM transition. As pointed out in Secs. II and V, FM-like RRs are formed around that transition. This is because a local FM coupling $J$ can be one of the strongest energy scale in the system [decimation case (iv) in Sec. V]. When this happens, the effective dissipation strength $\tilde{\alpha}$ increases with the volume of the RR [see Eqs. (27) and (28)]. Thus, a sufficiently large RR (and they can be arbitrarily large in a Griffiths phase (fisher_2)) have an effective dissipation parameter $\tilde{\alpha}>\alpha_{c}=1/2$. As a consequence, these larger RRs will cease tunneling ($\tilde{g}\rightarrow 0$) and undergo their own FM transition when Ohmic dissipation is considered. Thus, true local magnetic order develops. As this transition is not a collective effect of the bulk, the transition is said to be smeared. Any residual interaction between these larger RRs will align them together and the total magnetization will not average out. The Griffiths PROD phase of the undamped system now becomes a inhomogeneous FM phase due to Ohmic dissipation. Upon lowering $\delta$, this inhomogeneous FM phase crosses over to a weakly ordered phase (the former Griffiths FM phase of the undamped system) and, finally, crosses over to the conventional FM phase when further lowering $\delta.$ It is important to notice that this phenomenon is akin to the smearing in the PM-FM transition in the random transverse-field Ising chain studied in Refs. (smeared_hoyos_1; smeared_hoyos_2). Thus, both PM-FM ($\epsilon_{I}<1$) and PROD-FM ($\epsilon_{I}>1$) transitions become smeared in the presence of Ohmic dissipation as sketched in Fig. 1(b). We expect that all these conclusions remain valid for sub-Ohmic dissipation as well. For weak super-Ohmic dissipation, on the other hand, all transitions remain sharp because a RR cannot stop tunneling due to the coupling to the oscillator bath regardless the effective value of $\tilde{\alpha}$.
@@
It is interesting to quantify some key observables in those transitions. For the sharp transitions, the universality class is that of the random transverse-field Ising model which is extensively studied. All critical exponents are universal (i.e., independent of the details of the distributions of the coupling constants and fields) and well known. We refer the reader to Ref. (2014_ashkinteller) for the critical behavior of many observables. For the smearing transitions, many observables were quantified in Refs. (smeared_hoyos_1; smeared_hoyos_2). Here, we only quote the magnetization at the onset of the inhomogeneous FM phase: $\overline{\left\langle\sigma_{i}^{z}\right\rangle}=\overline{\left\langle S_{1,i}^{z}\right\rangle}=\overline{\left\langle S_{2,i}^{z}\right\rangle}\sim w^{1/\alpha},$ where $\overline{\left\langle\cdot\right\rangle}$ denotes the disorder and quantum averages. Here, $\alpha$ is the bare dissipation strength and $w$ is the probability of finding a local coupling constant $J_{i}$ being greater than $\gamma_{i}=\max\left\{h_{i}^{*},g_{i}^{*}\right\}$, where $h^{*}=h\left(h/\Omega_{I}\right)^{\frac{\alpha}{1-\alpha}}\approx h$ and, likewise, $g^{*}\approx g$. Precisely, $w=\int_{0}^{\infty}dJP_{J}(J)\int_{0}^{J}d\gamma P_{\gamma}\left(\gamma\right)$. Notice that $w$ plays the role of distance from transition.
VII Discussion and conclusions

### v2

@@ V Strong-disorder renormalization group
We now present the SDRG decimation procedure relevant for the $\epsilon_{I}>1$ regime in the presence of dissipation. We follow the methods of the dissipationless case in Ref. [30] and include dissipation using the strategy of Ref. [6] which is the following. The renormalization-group decimations are of two types: global and local, and are performed simultaneously when lowering the cutoff energy scale of the system. The global decimation concerns the integration of the fastest bath oscillators and is related to the adiabatic renormalization (see Sec. IV). The local decimation concerns the coarse grainning of a high-energy localized mode in a disordered system and, thus, is identical to the dissipationless case.
@@ V.1 Global decimation
where $\Omega_{I}$ is the bare $\Omega$. Notice that this is the global decimation procedure mentioned above. In contrast, the remaining steps are local decimations and are identical to the dissipationless case, except when it comes to renormalizing the dissipation parameter $\alpha_{i}$.
V.2 Local decimations
An interesting feature of the local decimations in the Ashkin-Teller model, when compared to the Ising model, is that there are two coupled dynamical variables in the problem: $\sigma$ and $\eta$ [see Eq. (4)]. At zeroth order, they are not locally coupled and, therefore, the decimation must take into account two local energy scales: one for each variable. This is clear when one considers two decoupled Ising models. A local decimation must take into account the largest energy scale in each model.
(ii) Let the largest energy scale in the system be $g_{i}$ and the second largest in that local cluster be $K_{i}$ (or vice versa). Thus the unperturbed Hamiltonian is $H_{0}=-K_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}-g{}_{i}\eta_{i}^{x}$ and all other terms in the Hamiltonian (4) containing $\boldsymbol{\sigma}_{i}$, $\boldsymbol{\sigma}_{i+1}$, and $\boldsymbol{\eta}_{i}$ are treated perturbatively up to second order. As a result, the new effective Hamiltonian has the same form as the Hamiltonian (4) but with one fewer site and two fewer spin-1/2 degrees of freedom. The degrees of freedom $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\sigma}_{i+1}$ are clustered together and replaced by an effective spin-1/2 degree of freedom $\tilde{\boldsymbol{\sigma}}$ upon which the effective field is $\tilde{h}$. In addition, $\boldsymbol{\eta}_{i}$ becomes locked in the $x$ direction and does not contribute to the $z$ magnetization. The effective coupling between the neighboring $\boldsymbol{\eta}$’s is $\tilde{J}$. This is sketched in Fig. 2(a), and the renormalized couplings are
$\tilde{J}=\frac{2J_{i-1}J_{i}}{\tilde{g}_{i}}\mbox{ and }\tilde{h}=\frac{2h_{i}h_{i+1}}{K_{i}}.$ (21)
Now we need to compute how the dissipation strength $\alpha$ changes under this decimation. We recall that
@@ V.2 Local decimations
and that the doubly degenerate ground state is $\left\{\left|\rightarrow_{i};\uparrow_{i}\uparrow_{i+1}\right\rangle,\left|\rightarrow_{i};\downarrow_{i}\downarrow_{i+1}\right\rangle\right\}$. Then, projecting $H_{\text{SB}}$ onto these states, we find that
@@
This is simply because $\left\langle\eta_{i}^{z}\right\rangle=\left\langle\eta_{i}^{z}\sigma_{i}^{z}\right\rangle=0$ and $\left\langle\sigma_{i+1}^{z}\right\rangle=\tilde{\sigma}^{z}$. Actually, this could be obtained in another fashion. Since $g_{i}$ is the largest energy scale, it means that the associated degree of freedom is actually already (or nearly) decoupled from the associated bath via successive adiabatic renormalizations (see Sec. IV). Thus only the last term in (22) needs to be projected onto the ground-state manifold. Importantly, there is no change in the spectral function of the remaining baths and, therefore, $\alpha_{i+1}$ remains unchanged.
Figure 2: Decimation procedure: (a) Global decimation due to the coupling with the dissipative bath [with $\tilde{h}_{i}$ and $\tilde{g}_{i}$ given by (19) and (20), respectively]. (b) Local decimations due to a broad distribution of local energy scales. When the two largest local energy scales in the system are (i) $g_{2}$ and $K_{2}$ [with $\tilde{J}$ and $\tilde{h}$ given by (21)], (ii) $g_{2}$ and $h_{2}$ [with $\tilde{K}$ and $\tilde{J}$ given by (24)], and (iii) $K_{2}$ and $J_{2}$ [with $\tilde{g}$ and $\tilde{h}$ given by (25)] In the main text, these correspond to the decimation cases (i), (ii), (iii) and (iv), respectively. The couplings between the spin clusters and the oscillator baths are also renormalized and are not illustrated here.
(iii) Let the largest energy scales be $g_{i}$ and the second largest in that local cluster be $h_{i}$. In this case, all tunneling frequencies have fully (or nearly) converged to their fixed-point values $g^{*}$ and $h^{*}$ in the adiabatic renormalization. The SDRG procedure is, thus, the same as the dissipationless case. The unperturbed Hamiltonian to be considered is $H_{0}=-h_{i}\sigma_{i}^{x}\left(1+\eta_{i}^{x}\right)-g_{i}\eta_{i}^{x}$. Projecting the spins $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\eta}_{i}$ onto the ground state of $H_{0}$ means that both spins are going to be decimated. The renormalized system is that illustrated in Fig. 2(b) and the renormalized quantities are
$\tilde{K}=\frac{K_{i-1}K_{i}}{2h_{i}}\mbox{ and }\tilde{J}=\frac{J_{i-1}J_{i}}{g_{i}+h_{i}}.$ (24)
@@
(iv) The final case to consider is when $K_{i}$ is the largest energy scale in the system and $J_{i}$ is the second largest in that local cluster. The unperturbed Hamiltonian is $H_{0}=-K_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}-J_{i}\eta_{i}^{z}\eta_{i+1}^{z}-J_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}\eta_{i}^{z}\eta_{i+1}^{z}$. Projecting $\boldsymbol{\sigma}_{i}$, $\boldsymbol{\sigma}_{i+1}$, $\boldsymbol{\eta}_{i}$, and $\boldsymbol{\eta}_{i+1}$ onto the ground state means that those spins are always aligned and, therefore, can be represented by two effective degrees of freedom $\tilde{\boldsymbol{\sigma}}$ and $\tilde{\boldsymbol{\eta}}$. Fusing them into clusters is illustrated in Fig. 2(c) where the renormalized quantities are
$\tilde{g}=\frac{g_{i}g_{i+1}}{2J_{i}}\mbox{ and }\tilde{h}=\frac{h_{i}h_{i+1}}{K_{i}+J_{i}}.$ (25)
For this step, we need to take a closer look at what happens to the oscillator baths. The system-bath Hamiltonian is
@@
and needs to be projected onto the fourfold-degenerate ground state of $H_{0}$, which is $\left\{\left|\eta;\sigma\right\rangle\right\}$ (where $\eta_{j}^{z}\left|\eta;\sigma\right\rangle=\eta\left|\eta;\sigma\right\rangle$, $\sigma_{j}^{z}\left|\eta;\sigma\right\rangle=\sigma\left|\eta;\sigma\right\rangle$, with $j=i,i+1$ and $\eta,\sigma=\pm 1$ representing the effective spins). Since $H_{\text{SB}}$ is diagonal in all four states, the projection can be easily performed, yielding the following effective Hamiltonian
@@
where the sum in the index $j$ runs from $i$ to $i+1$. Consequently, the spectral density of the local renormalized bath is
@@
It is thus interesting to notice that $\tilde{\alpha}$, increases under the RG flow only when FM-like clusterings take place, i.e., only when $J_{i}$ is decimated. Moreover, from Eq. (29), it becomes clear that the effective dissipation strength increases as the volume of the FM cluster, just like in the case of a direct transition between the PM and FM phases studied in Refs. [5, 6]. In addition, it is possible to relate the dissipation strength with the effective magnetic moment as both increase with the cluster volume. It is a convenient relation when analyzing the SDRG flow equations for the distributions of couplings, fields and dissipation parameters [6].
VI Phase diagram: dissipative case
@@
We are now able to describe the phase diagram of our model Hamiltonian (5) sketched in Fig. 1. A phase diagram is obtained after analyzing the corresponding strong-disorder renormalization-group flow of the Hamiltonian parameters. A careful reader may note that the decimation cases (i)–(iv) described in Sec. V do not cover all possible decimations. However, they do cover all decimations necessary to analyze the PM-PROD and PROD-FM transitions as shown in Ref. [30]. The missing decimations are those where $J_{i}$ or $h_{i}$ is the largest energy scale. They are important in analyzing the $\epsilon<1$ case where the two chains decouple and $\epsilon$ renormalizes to zero [29]. This situation, thus, was analyzed elsewhere with and without dissipation [29, 5, 6]
VI.1 Reviewing the dissipationless case
To briefly review the undamped case (see further details in Sec. II.3), we start noticing that the PM-PROD transition happens from the competition between the transverse field $h$ and the 4-spin coupling $K$ [see the first sum in the Hamiltonian (4); the other terms are just bystanders]. The relevant decimations are, thus, those described in cases (ii) and (iii) in the previous section. Comparing the corresponding renormalized values $\tilde{h}$ and $\tilde{K}$ in Eqs. (21) and (24), the transition line happens when the typical values are $2h_{\text{typ}}=K_{\text{typ}}$. Thus the transition line in the undamped case is at $\delta\approx\epsilon_{I}/2$. For the PROD-FM transition, the competition is between the 2-spin couplings $J$ and $g$ [see the second sum in the Hamiltonian (4)]. The relevant decimations are thus the cases (ii) and (iv). Following the same reasoning as before, the transition line is at $\delta\approx 2/\epsilon_{I}.$
VI.2 Dissipative case: Sharp transition
@@
Let us now analyze the effects of weak Ohmic dissipation $\alpha\ll 1$ and $s=1$ in Eq. (8)] in the PM-PROD transition. Since $\alpha\ll 1$ and does not renormalize, the decimation of fast oscillators has little effect on the renormalized values of $h$’s [see Eq. (19)]. Therefore, Ohmic dissipation is an irrelevant perturbation to the PM-PROD quantum phase transition. Its only effect is a slight change of the critical line. This same conclusion applies to the case of super-Ohmic dissipation ($s>1$) which is qualitatively weaker when compared to Ohmic dissipation [33]. Very likely, the same conclusion may also apply to the case of sub-Ohmic dissipation ($s<1$).
VI.3 Dissipative case: Smeared transition
@@
The effects of dissipation are much more dramatic in the PROD-FM transition. As pointed out in Secs. II and V, FM-like RRs are formed around that transition. This is because a local FM coupling $J$ can be the second largest energy scale in the local cluster [decimation case (iv) in Sec. V]. When this happens, the effective dissipation strength $\tilde{\alpha}$ increases with the volume of the RR [see Eqs. (28) and (29)]. Thus, a sufficiently large RR (and they can be arbitrarily large in a Griffiths phase [2]) has an effective dissipation parameter $\tilde{\alpha}>\alpha_{c}=1/2$. As a consequence, these larger RRs will cease tunneling ($\tilde{g}\rightarrow 0$) and undergo their own FM transition when Ohmic dissipation is considered. Thus, true local magnetic order develops. As this transition is not a collective effect of the bulk, the transition is said to be smeared. Any residual interaction between these larger RRs will align them together and the total magnetization will not average out. The Griffiths PROD phase of the undamped system now becomes an inhomogeneous FM phase due to Ohmic dissipation. Upon lowering $\delta$, this inhomogeneous FM phase crosses over to a weakly ordered phase (the former Griffiths FM phase of the undamped system) and, finally, crosses over to the conventional FM phase when further lowering $\delta.$ Having the different names for the IFM and WO phases is to emphasize that in the former, the magnetization appears only in some frozen and large rare regions while in the latter, the magnetization happens in the bulk. It is important to notice that this phenomenon is akin to the smearing in the PM-FM transition in the random transverse-field Ising chain studied in Refs. [5, 6]. Thus, both PM-FM ($\epsilon_{I}<1$) and PROD-FM ($\epsilon_{I}>1$) transitions become smeared in the presence of Ohmic dissipation as sketched in Fig. 1(b). We expect that all these conclusions remain valid for sub-Ohmic dissipation as well. For weak super-Ohmic dissipation, on the other hand, all transitions remain sharp because a RR cannot stop tunneling due to the coupling to the oscillator bath regardless the effective value of $\tilde{\alpha}$.
VI.4 Observables
@@
It is interesting to quantify some key observables in those transitions. For the sharp transitions, the universality class is that of the random transverse-field Ising model which is extensively studied. All critical exponents are universal (i.e., independent of the details of the distributions of the coupling constants and fields) and well known. We refer the reader to Ref. [30] for the critical behavior of many observables. For the smearing transitions, many observables were quantified in Refs. [5, 6]. Here, we only quote the magnetization at the onset of the inhomogeneous FM phase: $\overline{\left\langle\sigma_{i}^{z}\right\rangle}=\overline{\left\langle S_{1,i}^{z}\right\rangle}=\overline{\left\langle S_{2,i}^{z}\right\rangle}\sim w^{1/\alpha},$ where $\overline{\left\langle\cdot\right\rangle}$ denotes the disorder and quantum averages. Here, $\alpha$ is the bare dissipation strength and $w$ is the probability of finding a local coupling constant $J_{i}$ being greater than $\gamma_{i}=\max\left\{h_{i}^{*},g_{i}^{*}\right\}$, where $h^{*}=h\left(h/\Omega_{I}\right)^{\frac{\alpha}{1-\alpha}}\approx h$ and, likewise, $g^{*}\approx g$. Precisely, $w=\int_{0}^{\infty}dJP_{J}(J)\int_{0}^{J}d\gamma P_{\gamma}\left(\gamma\right)$. Notice that $w$ plays the role of distance from transition.
VII Discussions and conclusions